The following elements of shareholder’s equity are adapted from the balance sheet of Green Manufacturing Ltd.:
Shareholder’s Equity
Preferred shares, $0.75 cumulative,
100,000 shares authorized, 10,000 shares issued and outstanding $150,000


Common shares, 200,000 shares authorized, 60,000 shares issued
and outstanding 400,000
The company has paid all dividends through 2015
Compute the dividends paid to preferred shareholders and to common shareholders for 2018 and 2019 if total dividends declared are $40,000 in 2018 and $25,000 in 2019. Round your answers to nearest dollar.
Answer #1

Preferred dividend = 10000*.75 = 7500

Total Preferred Common
2018 40000 7500*3 = 22500 40000-22500 = 17500
2019 25000 7500 25000-7500 = 17500